# Learning 3D Representations from Procedural 3D Programs

### v2

@@ 4 Experiments
We present a comprehensive evaluation of 3D shape representations pretrained with procedural 3D programs across various downstream object-level and scene-level tasks, including object classification, part segmentation, and 3D scene understanding (Sec. 4.1 – 4.4). We further provide an in-depth analysis of model behavior and ablation studies (Sec. 4.5). For each downstream task, we report the performance of relevant existing methods as a reference and focus on comparisons with SSLs pretrained on manually-curated 3D datasets, as well as models trained from scratch. Specifically, for object-level 3D understanding tasks, we evaluate the following three pretraining strategies: (1) Scratch: All network parameters are randomly initialized, with no pretraining. (2) ShapeNet Pretrained (SN): Pretrained on 41,952 models in the ShapeNet [4] training split, relying on the officially released weights. (3) Procedural 3D Programs Pretrained (Zero): Pretrained on 150K procedurally generated 3D models, using no human-crafted shapes. For scene-level tasks, we compare SSL models pretrained on ScanNet [7] and procedurally generated 3D scenes.
@@ 4.1 Object Classification
Methods ModelNet40 OBJ-BG OBJ-ONLY PB-T50-RS
PointNet [27] 89.2 73.3 79.2 68.0
SpiderCNN [48] 92.4 77.1 79.5 73.7
PointNet++ [28] 90.7 82.3 84.3 77.9
DGCNN [38] 92.9 86.1 85.5 78.5
PointCNN [19] 86.1 85.5 78.5
PTv1 [62] 93.7
PTv2 [40] 94.2
OcCo [37] 92.1 84.9 85.5 78.8
Point-BERT [55] 93.2 87.4 88.1 83.1
Point-MAE-Scratch [55] 91.4 79.9 80.6 77.2
Point-MAE-SN [26] 93.8 90.0 88.3 85.2
Point-MAE-Zero 93.0 90.4 88.6 85.5
Point-M2AE-Scratch 92.2 90.0 87.6 85.6
Point-M2AE-SN [59] 94.0 91.2 88.8 86.4
Point-M2AE-Zero 92.9 90.4 89.8 87.0
PCP-MAE-Scratch 91.5 88.8 88.5 83.8
PCP-MAE-SN [61] 94.0 95.5 94.3 90.4
PCP-MAE-Zero 92.4 94.0 92.3 90.5
Table 1: Object Classification. We evaluate the object classification performance on ModelNet40 and three variants of ScanObjectNN. Classification accuracy (%) is reported (higher is better). Top: Performance of existing methods with various neural network architectures and pretraining strategies. Bottom: Comparison with our baseline methods.
@@
We use ModelNet40 [44] and ScanObjectNN [34] as the benchmarks for the shape classification task. ModelNet40 contains 12,311 clean 3D CAD objects across 40 categories, with 9,843 samples for training and 2,468 for testing. Following Point-MAE, we apply random scaling and translation as data augmentation during training, and a voting strategy during testing [22]. Following prior works [37, 55, 26], we also evaluate the few-shot classification performance on ModelNet40. ScanObjectNN is a more complex real-world 3D dataset, consisting of approximately 15,000 objects across 15 categories, with items scanned from cluttered indoor scenes. We report results on three ScanObjectNN variants: OBJ-BG, OBJ-ONLY, and PB-T50-RS, the latter being the most challenging due to its additional noise and occlusions.
@@
Table 1 summarizes object classification results across several settings. On ModelNet40, the “-Zero” variants (e.g., Point-MAE-Zero, Point-M2AE-Zero, PCP-MAE-Zero) generally fall slightly behind their ShapeNet‐pretrained counterparts (“-SN”), reflecting the larger domain gap between synthetic shapes and the clean 3D models in ModelNet40. By contrast, on ScanObjectNN—which contains real-world scans with broader geometric variability—the “-Zero” models often match or exceed the performance of their “-SN” counterparts. For instance, PCP-MAE-Zero outperforms PCP-MAE-SN on the PB-T50-RS variant, and Point-M2AE-Zero closely matches or exceeds Point-M2AE-SN in several cases. These findings indicate that the diverse geometry in procedurally synthesized data can be advantageous for certain real-world tasks. Meanwhile, all pretrained models (including both “-SN” and “-Zero”) surpass their respective from-scratch baselines and outperform existing self-supervised approaches (e.g., OcCo [37] and Point-BERT [55]).
Methods 5w/10s 5w/20s 10w/10s 10w/20s
DGCNN-rand [38] 31.6$\pm$2.8 40.8$\pm$4.6 19.9$\pm$2.1 16.9$\pm$1.5
DGCNN-OcCo [38] 90.6$\pm$2.8 92.5$\pm$1.9 82.9$\pm$1.3 86.5$\pm$2.2
Transformer-OcCo [37] 94.0$\pm$3.6 95.9$\pm$2.3 89.4$\pm$5.1 92.4$\pm$4.6
Point-BERT [55] 94.6$\pm$3.1 96.3$\pm$2.7 91.0$\pm$5.4 92.7$\pm$5.1
Point-MAE-Scratch [55] 87.8$\pm$5.2 93.3$\pm$4.3 84.6$\pm$5.5 89.4$\pm$6.3
Point-MAE-SN [26] 96.3$\pm$2.5 97.8$\pm$1.8 92.6$\pm$4.1 95.0$\pm$3.0
Point-MAE-Zero 95.4$\pm$2.5 97.7$\pm$1.6 91.3$\pm$5.1 95.0$\pm$3.5
Point-M2AE-Scratch 87.5$\pm$2.6 90.0$\pm$5.5 86.4$\pm$3.2 89.6$\pm$4.3
Point-M2AE-SN [26] 93.4$\pm$3.1 96.2$\pm$1.5 91.8$\pm$4.5 92.9±$\pm 3.2$
Point-M2AE-Zero 91.4$\pm$1.8 94.2$\pm$2.2 88.3$\pm$3.7 91.0$\pm$2.9
PCP-MAE-Scratch 86.4$\pm$2.6 85.0$\pm$6.0 88.9$\pm$4.1 90.7$\pm$4.2
PCP-MAE-SN [61] 97.4$\pm$2.3 99.1$\pm$0.8 93.5$\pm$3.7 95.9$\pm$2.7
PCP-MAE-Zero 95.3$\pm$3.4 98.4$\pm$1.4 91.5$\pm$4.4 94.7$\pm$3.0
Table 2: Few-shot classification on ModelNet40. We evaluate performance on four $n$-way, $m$-shot configurations. For example, 5w/10s denotes a 5-way, 10-shot classification task. The table reports the mean classification accuracy (%) and standard deviation across 10 independent runs for each configuration. Top: Results from existing methods for comparison. Bottom: Comparison with our baseline methods. Note that results for Point-M2AE-SN are reproduced using publicly available code with our own configuration, as the original configuration was not provided.
@@
We evaluate few-shot classification on ModelNet40 using standard $n$-way, $m$-shot protocols, where $n$ denotes the number of randomly selected classes and $m$ the number of examples per class. Each evaluation samples 20 unseen instances from each class. We repeat this procedure 10 times, reporting mean accuracy (%) and standard deviation. Table 2 presents results for $n=\{5,10\}$ and $m=\{10,20\}$. Similar to transfer learning experiments, Point-MAE-Zero performs slightly below Point-MAE-SN, likely due to the larger domain gap between procedural shapes and ModelNet40 data. Nevertheless, both methods significantly outperform scratch-trained models and previous approaches such as DGCNN [38] and Transformer-OcCo [37].
@@ 4.2 Part Segmentation
The 3D part segmentation task aims to assign a part label to each point in a shape. We evaluate our methods and baselines on ShapeNetPart [54], which contains 16,881 models across 16 object categories. Consistent with previous works [27, 26, 55], we sample 2,048 points from each shape, resulting in 128 patches in our masked autoencoding pipeline (see Sec. 3).
@@
In line with our observations in Sec. 4.1, the “-Zero” and “-SN” models surpass scratch-trained baselines and earlier methods that use different architectures [27, 28, 38] or alternative pretraining strategies [37, 55]. Despite lacking high-level semantic cues, these procedurally trained autoencoders still capture sufficient geometric structure to achieve strong segmentation performance.
@@
PointNet [27] 83.7 83.4 78.7 82.5 74.9 89.6 73.0 91.5 85.9
PointNet++ [28] 85.1 82.4 79.0 87.7 77.3 90.8 71.8 91.0 85.9
DGCNN [38] 85.2 84.0 83.4 86.7 77.8 90.6 74.7 91.2 87.5
OcCo [37] 85.1 83.3 85.2 88.3 79.9 90.7 74.1 91.9 87.6
Point-BERT [55] 85.6 84.3 84.8 88.0 79.8 91.0 81.7 91.6 87.9
Point-MAE-Scratch [55] 85.1 82.9 85.4 87.7 78.8 90.5 80.8 91.1 87.7
Point-MAE-SN [26] 86.1 84.3 85.0 88.3 80.5 91.3 78.5 92.1 87.4
Point-MAE-Zero 86.1 85.0 84.2 88.9 81.5 91.6 76.9 92.1 87.6
@@
Point-M2AE-SN [59] 85.0 84.5 87.2 89.3 81.1 91.8 80.1 92.0 89.2
Point-M2AE-Zero 84.9 85.3 87.3 88.7 81.1 91.7 79.4 91.9 88.2
@@
PointNet [27] 80.8 95.3 65.2 93.0 81.2 57.9 72.8 80.6
PointNet++ [28] 83.7 95.3 71.6 94.1 81.3 58.7 76.4 82.6
DGCNN [38] 82.8 95.7 66.3 94.9 81.1 63.5 74.5 82.6
OcCo [37] 84.7 95.4 75.5 94.4 84.1 63.1 75.7 80.8
Point-BERT [55] 85.2 95.6 75.6 94.7 84.3 63.4 76.3 81.5
Point-MAE-Scratch [55] 85.3 95.6 73.9 94.9 83.5 61.2 74.9 80.6
Point-MAE-SN [26] 86.1 96.1 75.2 94.6 84.7 63.5 77.1 82.4
@@
Table 3: Part Segmentation Results. We report the mean Intersection over Union (IoU) across all instances (mIoUI) and the IoU (%) for each category on the ShapeNetPart benchmark (higher values indicate better performance).
Figure 3: Masked Point Cloud Completion. This figure visualizes shape completion results with Point-MAE-SN and Point-MAE-Zero on the ShapeNet test split and procedurally synthesized 3D shapes. Left: Ground truth point clouds and masked inputs (60% mask ratio). Middle: Completions guided by masked input patch centers [26]. Right: Reconstructions without any guidance points. The $L_{2}$ Chamfer distance (lower is better) between the predicted 3D point clouds and the ground truth is displayed below each reconstruction.
@@ 4.3 Masked Point Cloud Completion
The masked point cloud completion task aims to reconstruct masked regions of input 3D point clouds, serving as a self-supervised pretext for learning 3D representations [26] (see Fig. 2 and Sec. 3).
@@
Semantic Segmentation Instance Segmentation
Methods mIoU mAcc allAcc mAP AP50 AP25
MSC-scan [41] 73.85 81.80 90.49 39.75 60.51 76.49
MSC-Zero (1k) 72.69 80.80 90.13 39.03 58.57 75.24
MSC-Zero (2k) 73.86 82.03 91.18 40.81 62.28 76.26
@@
4.4 Scene-level 3D Understanding Tasks
Given the effectiveness of procedural 3D programs for pretraining self-supervised learning (SSL) models on 3D objects, a natural question arises: Can procedural 3D programs similarly benefit SSL for 3D scenes? We adopt Masked Scene Contrast (MSC) [41], a state-of-the-art SSL method for 3D scenes. We pretrain MSC on ScanNet [7] (1K scenes), commonly used for 3D scene SSL pretraining, and compare it against MSC pretrained on our procedurally generated scenes (denoted MSC-Zero). We conduct experiments with MSC-Zero using varying amounts of data (1K and 2K procedural scenes). MSC-Zero trained with 2K procedurally generated 3D scenes achieves outperforms MSC pretrained on ScanNet in both 3D semantic and instance segmentation tasks. This aligns with our observations on object-level 3D understanding tasks. We discuss the exact procedures of generating such data and implementation details in the supplementary materials.
@@ 4.4 Scene-level 3D Understanding Tasks
Figure 4: Impact of 3D Shape Complexity on Performance. Left: Examples of procedurally generated 3D shapes with increasing complexity, used for pretraining. Textures are shown for illustration purposes only; in practice, only the surface points are used. Right: Comparison of pretraining masked point reconstruction loss (Eqn. 1[26] and downstream classification accuracy on the ScanObjectNN dataset [34]. Each row in Point-MAE-Zero represents an incrementally compounded effect of increasing shape complexity and augmentation, with the highest accuracy achieved using shape augmentation.

### v3

@@ 4 Experiments
We present a comprehensive evaluation of 3D shape representations pretrained with procedural 3D programs across various downstream object-level and scene-level tasks, including object classification, part segmentation, and 3D scene understanding (Sec. 4.1 – 4.4). We further provide an in-depth analysis of model behavior and ablation studies (Sec. 4.5). For each downstream task, we report the performance of relevant existing methods as a reference and focus on comparisons with SSLs pretrained on manually-curated 3D datasets, as well as models trained from scratch. Specifically, for object-level 3D understanding tasks, we evaluate the following three pretraining strategies: (1) Scratch: All network parameters are randomly initialized, with no pretraining. (2) ShapeNet Pretrained (SN): Pretrained on 41,952 models in the ShapeNet [chang2015shapenet] training split, relying on the officially released weights. (3) Procedural 3D Programs Pretrained (Zero): Pretrained on 150K procedurally generated 3D models, using no human-crafted shapes. For scene-level tasks, we compare SSL models pretrained on ScanNet [dai2017scannet] and procedurally generated 3D scenes.
@@ 4.1 Object Classification
Methods ModelNet40 OBJ-BG OBJ-ONLY PB-T50-RS Avg.
PointNet [qi2017pointnet] 89.2 73.3 79.2 68.0 77.4
SpiderCNN [xu2018spidercnn] 92.4 77.1 79.5 73.7 80.7
PointNet++ [qi2017pointnet++] 90.7 82.3 84.3 77.9 83.8
DGCNN [wang2019dynamic] 92.9 86.1 85.5 78.5 85.8
PointCNN [li2018pointcnn] 86.1 85.5 78.5
PTv1 [zhao2021point] 93.7
PTv2 [wu2022point] 94.2
OcCo [wang2021occo] 92.1 84.9 85.5 78.8 85.3
Point-BERT [yu2022point] 93.2 87.4 88.1 83.1 88.0
Point-MAE-Scratch [yu2022point] 91.4 79.9 80.6 77.2 82.3
Point-MAE-SN [pang2022masked] 93.8 90.0 88.3 85.2 89.3 (+7.0)
Point-MAE-Zero 93.0 90.4 88.6 85.5 89.4 (+7.1)
Point-M2AE-Scratch 92.2 90.0 87.6 85.6 88.9
Point-M2AE-SN [zhang2022point] 94.0 91.2 88.8 86.4 90.1 (+1.2)
Point-M2AE-Zero 92.9 90.4 89.8 87.0 90.0 (+1.1)
PCP-MAE-Scratch 91.5 88.8 88.5 83.8 88.2
PCP-MAE-SN [zhang2024pcp] 94.0 95.5 94.3 90.4 93.6 (+5.4)
PCP-MAE-Zero 92.4 94.0 92.3 90.5 92.3 (+4.1)
Table 1: Object Classification. We evaluate the object classification performance on ModelNet40 and three variants of ScanObjectNN. Classification accuracy (%) is reported (higher is better). Top: Performance of existing methods with various neural network architectures and pretraining strategies. Bottom: Comparison with our baseline methods. The rightmost column shows the average accuracy, with red text indicating the improvement over the corresponding Scratch baseline.
@@
We use ModelNet40 [wu20153d] and ScanObjectNN [uy2019revisiting] as the benchmarks for the shape classification task. ModelNet40 contains 12,311 clean 3D CAD objects across 40 categories, with 9,843 samples for training and 2,468 for testing. Following Point-MAE, we apply random scaling and translation as data augmentation during training, and a voting strategy during testing [liu2019relation]. Following prior works [wang2021occo, yu2022point, pang2022masked], we also evaluate the few-shot classification performance on ModelNet40. ScanObjectNN is a more complex real-world 3D dataset, consisting of approximately 15,000 objects across 15 categories, with items scanned from cluttered indoor scenes. We report results on three ScanObjectNN variants: OBJ-BG, OBJ-ONLY, and PB-T50-RS, the latter being the most challenging due to its additional noise and occlusions.
@@
Table 1 summarizes object classification results across several settings. On ModelNet40, the “-Zero” variants (e.g., Point-MAE-Zero, Point-M2AE-Zero, PCP-MAE-Zero) generally fall slightly behind their ShapeNet‐pretrained counterparts (“-SN”), reflecting the larger domain gap between synthetic shapes and the clean 3D models in ModelNet40. By contrast, on ScanObjectNN—which contains real-world scans with broader geometric variability—the “-Zero” models often match or exceed the performance of their “-SN” counterparts. For instance, PCP-MAE-Zero outperforms PCP-MAE-SN on the PB-T50-RS variant, and Point-M2AE-Zero closely matches or exceeds Point-M2AE-SN in several cases. These findings indicate that the diverse geometry in procedurally synthesized data can be advantageous for certain real-world tasks. Meanwhile, all pretrained models (including both “-SN” and “-Zero”) surpass their respective from-scratch baselines and outperform existing self-supervised approaches [wang2021occo, yu2022point] which we highlight with the rightmost column.
Methods 5w/10s 5w/20s 10w/10s 10w/20s Avg.
[…]
Table 2: Few-shot classification on ModelNet40. We evaluate performance on four $n$-way, $m$-shot configurations. For example, 5w/10s denotes a 5-way, 10-shot classification task. The table reports the mean classification accuracy (%) and standard deviation across 10 independent runs for each configuration. Top: Results from existing methods for comparison. Bottom: Comparison with our baseline methods. Note that results for Point-M2AE-SN are reproduced using publicly available code, as the original configuration was not provided. The final column shows the average accuracy across configurations, with subscripts indicating improvements over the corresponding baseline.
@@
We evaluate few-shot classification on ModelNet40 using standard $n$-way, $m$-shot protocols, where $n$ denotes the number of randomly selected classes and $m$ the number of examples per class. Each evaluation samples 20 unseen instances from each class. We repeat this procedure 10 times, reporting mean accuracy (%) and standard deviation. Table 2 presents results for $n=\{5,10\}$ and $m=\{10,20\}$. Similar to transfer learning experiments, Point-MAE-Zero performs on par or slightly below Point-MAE-SN, likely due to the larger domain gap between procedural shapes and ModelNet40 data. Nonetheless, both methods substantially outperform their scratch-trained counterparts, as reflected by the performance deltas, and also surpass prior approaches such as DGCNN [wang2019dynamic] and Transformer-OcCo [wang2021occo].
@@ 4.2 Part Segmentation
The 3D part segmentation task aims to assign a part label to each point in a shape. We evaluate our methods and baselines on ShapeNetPart [yi2016scalable], which contains 16,881 models across 16 object categories. Consistent with previous works [qi2017pointnet, pang2022masked, yu2022point], we sample 2,048 points from each shape, resulting in 128 patches in our masked autoencoding pipeline (see Sec. 3).
@@
In line with our observations in Sec. 4.1, the “-Zero” and “-SN” models surpass scratch-trained baselines and earlier methods that use different architectures [qi2017pointnet, qi2017pointnet++, wang2019dynamic] or alternative pretraining strategies [wang2021occo, yu2022point]. Despite lacking high-level semantic cues, these procedurally trained autoencoders still capture sufficient geometric structure to achieve strong segmentation performance.
@@
PointNet [qi2017pointnet] 83.7 83.4 78.7 82.5 74.9 89.6 73.0 91.5 85.9
PointNet++ [qi2017pointnet++] 85.1 82.4 79.0 87.7 77.3 90.8 71.8 91.0 85.9
DGCNN [wang2019dynamic] 85.2 84.0 83.4 86.7 77.8 90.6 74.7 91.2 87.5
OcCo [wang2021occo] 85.1 83.3 85.2 88.3 79.9 90.7 74.1 91.9 87.6
Point-BERT [yu2022point] 85.6 84.3 84.8 88.0 79.8 91.0 81.7 91.6 87.9
Point-MAE-Scratch [yu2022point] 85.1 82.9 85.4 87.7 78.8 90.5 80.8 91.1 87.7
Point-MAE-SN [pang2022masked] 86.1(+1.0) 84.3 85.0 88.3 80.5 91.3 78.5 92.1 87.4
Point-MAE-Zero 86.1(+1.0) 85.0 84.2 88.9 81.5 91.6 76.9 92.1 87.6
@@
Point-M2AE-SN [zhang2022point] 85.0(+0.3) 84.5 87.2 89.3 81.1 91.8 80.1 92.0 89.2
Point-M2AE-Zero 84.9(+0.2) 85.3 87.3 88.7 81.1 91.7 79.4 91.9 88.2
@@
PointNet [qi2017pointnet] 80.8 95.3 65.2 93.0 81.2 57.9 72.8 80.6
PointNet++ [qi2017pointnet++] 83.7 95.3 71.6 94.1 81.3 58.7 76.4 82.6
DGCNN [wang2019dynamic] 82.8 95.7 66.3 94.9 81.1 63.5 74.5 82.6
OcCo [wang2021occo] 84.7 95.4 75.5 94.4 84.1 63.1 75.7 80.8
Point-BERT [yu2022point] 85.2 95.6 75.6 94.7 84.3 63.4 76.3 81.5
Point-MAE-Scratch [yu2022point] 85.3 95.6 73.9 94.9 83.5 61.2 74.9 80.6
Point-MAE-SN [pang2022masked] 86.1 96.1 75.2 94.6 84.7 63.5 77.1 82.4
@@
Table 3: Part Segmentation Results. We report the mean Intersection over Union over instances (mIoUI) and the per-category IoU (%) on the ShapeNetPart benchmark. Higher values indicate better performance. The subscripted values in mIoUI represent performance improvement over the corresponding baseline.
Figure 3: Masked Point Cloud Completion. This figure visualizes shape completion results with Point-MAE-SN and Point-MAE-Zero on the ShapeNet test split and procedurally synthesized 3D shapes. Left: Ground truth point clouds and masked inputs (60% mask ratio). Middle: Completions guided by masked input patch centers [pang2022masked]. Right: Reconstructions without any guidance points. The $L_{2}$ Chamfer distance (lower is better) between the predicted 3D point clouds and the ground truth is displayed below each reconstruction.
@@ 4.3 Masked Point Cloud Completion
Masked point cloud completion reconstructs missing regions of 3D point clouds as a self-supervised pretext task for learning 3D representations [pang2022masked] (see Fig. 2 and Sec. 3).
@@
4.4 Scene-level 3D Understanding Tasks
Semantic Seg. Instance Segmentation
Methods mIoU mAcc mAP AP50 AP25
MSC-scan [wu2023maskedscenecontrastscalable] 73.85 81.80 39.75 60.51 76.49
MSC-Zero (1k) 72.69 80.80 39.03 58.57 75.24
MSC-Zero (2k) 73.86 82.03 40.81 62.28 76.26
MSC-Zero (4k) 74.34 82.16 41.47 63.12 76.54
@@ 4.4 Scene-level 3D Understanding Tasks
Given the effectiveness of procedural 3D programs for pretraining self-supervised learning (SSL) models on 3D objects, a natural question arises: Can procedural 3D programs similarly benefit SSL for 3D scenes? We adopt Masked Scene Contrast (MSC) [wu2023maskedscenecontrastscalable], a popular SSL method for 3D scenes. We pretrain MSC on ScanNet [dai2017scannet] (1K scenes), commonly used for 3D scene SSL pretraining, and compare it against MSC pretrained on our procedurally generated scenes (denoted MSC-Zero). We conduct experiments with MSC-Zero using varying amounts of data (1K, 2K and 4K procedural scenes). MSC-Zero trained with 4K procedurally generated 3D scenes achieves outperforms MSC pretrained on ScanNet in both 3D semantic and instance segmentation tasks. This demonstrates the effectiveness of procedurally generated data for 3D scene self-supervised learning and reinforces our findings from object-level 3D understanding tasks. Moreover, our results demonstrate that increasing the number of procedural scenes consistently improves performance across semantic and instance segmentation tasks. We discuss the exact procedures of generating such data and implementation details in the supplementary materials.
@@
Figure 4: Impact of 3D Shape Complexity on Performance. Left: Examples of procedurally generated 3D shapes with increasing complexity, used for pretraining. Textures are shown for illustration purposes only; in practice, only the surface points are used. Right: Comparison of pretraining masked point reconstruction loss (Eqn. 1[pang2022masked] and downstream classification accuracy on the ScanObjectNN dataset [uy2019revisiting]. Each row in Point-MAE-Zero represents an incrementally compounded effect of increasing shape complexity and augmentation, with the highest accuracy achieved using shape augmentation.